a particle of mass m driven by amachine that delliver a constant power k watt . if the particle starts from rest the force of the particle at time t
 plzz give full solution
 

Grade:12th pass

3 Answers

akash
8 years ago
i think the question is incorrect. verify it.
 
Nithya
8 years ago
p=FV=K
K=mv dv/dt
integrate,
v2/2=kt/m
find v
a=dv/dt
F= ma= (mk/2t) power half
Abhishek Chaudhary
4 years ago
P=fv= mav                                                                         Now put a= dv/dt and integrate it to get value of v. When you get value of v then differentiate to get value of a